OpenStudy (anonymous):

Has anyone taken Algebra 2 yet?

OpenStudy (solomonzelman):

I took trig....

OpenStudy (anonymous):

I need Algebra 2 help

OpenStudy (solomonzelman):

Ask a question.

OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? x^4-52x^2+576=0

OpenStudy (anonymous):

A.) 4, -4 B.) 4, -6 C.) 4, -4, 6, -6 D.) no solution

OpenStudy (solomonzelman):

This is a trigonometric question, But I'll help you with it anyway.

OpenStudy (solomonzelman):

Say the following, let x^2=a, substitute a for x^2 x^4-52x^2+576=0 ---> a^2-52a+576=0 can you solve for a?

OpenStudy (solomonzelman):

@Coop21, do you need help with this?

OpenStudy (anonymous):

Yeah, I'm lost

OpenStudy (solomonzelman):

Do you see how I got a^2-52a+576=0?

OpenStudy (anonymous):

I'm so far behind I don't even know how to do anything in this class.

OpenStudy (anonymous):

No I don't

OpenStudy (solomonzelman):

Ok, we are making x^2=a right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

So we just called x^2 'a' so lets substitute a instead of x^2, OK?

OpenStudy (anonymous):

Ther is no a in the equation

OpenStudy (solomonzelman):

Yes, but knowing that x^2=a we are going to substitute a for x^2, solve for a, and then substitute the result for a into a=x^2. Right?

OpenStudy (solomonzelman):

Are you still here, @Coop21?

OpenStudy (anonymous):

but it is x^4 not x^2

OpenStudy (solomonzelman):

yes, but we are saying that x^2=a right?

OpenStudy (solomonzelman):

So x^4=a^2 and x^2=a, right?

OpenStudy (anonymous):

Sure

OpenStudy (solomonzelman):

Ok, so now we can say that x^4-52x^2+576=0 ---> a^2-52a+576=0

OpenStudy (solomonzelman):

Right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Which methods do you know to solve this? Factoring Quadratic formula Completing the Square

OpenStudy (anonymous):

Completing the Square?

OpenStudy (solomonzelman):

If you don't know how to complete the square, do you want to skip it for now, or learn how to do it?

OpenStudy (anonymous):

I have no clue about any of this

OpenStudy (solomonzelman):

not even the quadratic formula?

OpenStudy (anonymous):

I taking it now and were doing solving Quadratic Equation by completing the square

OpenStudy (anonymous):

Sloving polynomial equations

OpenStudy (solomonzelman):

Cool, @Misswright224. Which one do you like better? Quadratic Equation, completing the square, or factoring? I like completing square, it rocks.

OpenStudy (anonymous):

Completing the square